The free-throw line in basketball is 4.57 m from the basket, which is 3.05 mabove the floor. A player standing on the free-throw line throws the ball with an initial speed of 7.15 m/s, releasing it at a height of 2.44 m above the floor. At what angle above the horizontal must the ball be thrown to exactly hit the basket?

Note that most players will use a large initial angle rather than a flat shot because itallows for a larger margin of error.


1
Expert's answer
2022-03-07T11:45:03-0500

"x=v_x t,"

"y=y_0+v_yt-\frac{gt^2}2,"

"1.54\tan^2\alpha-4.57\tan\alpha+2.15=0,"

"\alpha=67.2°."